OpenStudy (anonymous):

A pair of shoes usually sells for $61. If the shoes are 20% off, and sales tax is 5%, what is the total price of the shoes, including tax?

OpenStudy (anonymous):

Discount price equals original price minus twenty percent of the original price. 61-(61*0.2)= Discount price Discount price + sales tax = out-the-door price Sales tax is discount price times 0.05 Discount price + (discount price times 0.05) = out-the-door price